### v1

# CMIC: Content-Adaptive Mamba for Learned Image Compression

Abstract
Recent Learned image compression (LIC) leverages Mamba-style state-space models (SSMs) for global receptive fields with linear complexity. However, vanilla Mamba is content-agnostic, relying on fixed and predefined selective scans, which restricts its ability to dynamically and fully exploit content dependencies. We introduce Content-Adaptive Mamba (CAM), a dynamic SSM that addresses two critical limitations. First, it employs content-aware token reorganization, clustering and reordering tokens based on content similarity to prioritize proximity in feature space over Euclidean space. Second, it integrates global priors into SSM via a prompt dictionary, effectively mitigating the strict causality and long-range decay in the token interactions of Mamba. These innovations enable CAM to better capture global dependencies while preserving computational efficiency. Leveraging CAM, our Content-Adaptive Mamba-based LIC model (CMIC) achieves state-of-the-art rate-distortion performance, surpassing VTM-21.0 by -15.91%, -21.34%, and -17.58% BD-rate on Kodak, Tecnick, and CLIC benchmarks, respectively.
1 Introduction
Image compression plays a vital role in computer vision and multimedia by reducing storage and bandwidth while preserving visual quality. Learned image compression (LIC) has advanced rapidly, driven by variational autoencoders and end-to-end training (Ballé et al. (2017, 2018)). Early LIC methods used convolutional neural networks (CNNs) for nonlinear analysis and synthesis transforms (Minnen et al. (2018); Cheng et al. (2020); He et al. (2022)). More recently, transformer-based designs expanded receptive fields to capture long-range dependencies (Zou et al. (2022); Zhu et al. (2022); Liu et al. (2023); Li et al. (2024a)). However, transformers achieve a global receptive field at the cost of quadratic time and memory complexity. State space models (SSMs) such as Mamba (Gu and Dao (2023); Dao and Gu (2024); Zhu et al. ; Liu et al. (2024)) offer a compelling alternative: they couple global receptive fields with linear complexity and have already been adopted in several LIC models (Qin et al. (2024); Zeng et al. (2025)).
Yet Mamba’s Selective Scan mechanism, originally designed for one-dimensional and causal sequences, faces significant challenges when applied to the inherently two-dimensional nature of images. First, the long-range decay phenomenon causes the influence of distant tokens to diminish rapidly along the sequence Guo et al. (2024a). The problem is compounded by the fixed selective scan mechanism, which rigidly processes tokens in a predefined order without considering their content similarities. The rigid, content-agnostic approach only addresses Euclidean adjacency but fails to exploit the rich content dependencies inherent in images. Semantically neighboring regions often share similar or related visual patterns, and redundancy also typically originates from these areas. Ignoring these relationships, vanilla Mamba potentially scans similar tokens in a non-continuous manner. It weakens Mamba’s ability to capture dependencies between tokens that are far apart in Euclidean space but close in feature space, ultimately leading to suboptimal rate-distortion (RD) performance. Second, vanilla Mamba’s strict causality restricts each token to accessing only its predecessors in the sequence, leaving subsequent tokens unexplored Liu et al. (2024); Guo et al. (2024b), contradicting the non-causal, interconnected nature of 2D images. These limitations collectively reduce the model’s ability to capture global dependencies, which are critical for high-performance image compression. While multi-directional 2-D scans partially mitigate these issues, they quadruple computational complexity and still rely on Euclidean adjacency for scans (Fig. 1), failing to adapt to content similarities. These shortcomings highlight the need for a more flexible and global-aware approach to modeling image data.
To overcome these limitations, we introduce Content-Adaptive Mamba (CAM), a dynamic SSM tailored to images. First, we address the content-agnostic fixed scan by clustering and reorganizing latent tokens to group content-similar tokens together. This ensures that tokens are processed in sequence based on their feature-space proximity rather than their spatial (Euclidean) arrangement, encouraging interaction between content-similar tokens. Intuitively, in the context of image compression, the core of nonlinear transformation lies in identifying tokens with similar content and eliminating redundancy between them. By enabling Mamba to perform continuous scanning for similar tokens, we preserve linear complexity while enhancing long-range interactions between similar tokens.
Second, we mitigate the strict causality and long-range decay by introducing a prompt dictionary to enhance the model’s global awareness. Compact prototype vectors are learned to capture shared semantic information for each cluster, while instance-specific prompts are dynamically generated from the dictionary. These prompts are integrated into the SSM, modulating hidden states and relaxing the strict causality with minimal computational overhead. By embedding categorical information and token relationships into the SSM, we expand its context and dynamism, enabling it to capture global dependencies more effectively.
Together, these two strategies enable efficient, content-aware long-range modeling. Built on them, our Content-Adaptive Mamba-based LIC model, CMIC, attains state-of-the-art RD performance, surpassing the traditional codec VTM-21.0 by -15.91%, -21.34%, and -17.58% BD-rate on the Kodak, Tecnick, and CLIC datasets, respectively. Our main contributions can be summarized as:
Figure 1: (a) Illustration of standard 2D Selective Scan. (b) Illustration of our content-adaptive scan. (c) Rate ravings relative to VTM-21.0 on the Tecknick dataset. CMIC significantly outperforms two SOTA Mamba-based LIC models: MambaVC (Qin et al. (2024)) and MambaIC (Zeng et al. (2025)).
  • Content-Adaptive Token Aggregation. We propose content-aware token reorganization, clustering and reordering tokens by feature similarity to prioritize feature-space proximity over Euclidean adjacency for Mamba, which strengthens long-range interactions.
  • Prompt Dictionary Conditioning. We introduce a learnable prompt dictionary that supplies instance-specific global priors, relaxing the strict causality and enhancing global dependency modeling for SSM with minimal overhead.
  • End-to-End CMIC Model. We build CMIC, a dynamic Mamba-based LIC model that achieves state-of-the-art rate-distortion performance across all the three datasets.
2 Related Work
2.1 Learned Image Compression
End-to-end learned image compression (LIC) has rapidly progressed in recent years. The original CNN-based method by Ballé et al. (2017)-and its extension that couples VAEs with hyper-priors (Ballé et al. (2018))—laid the groundwork for modern LIC. Subsequent studies have mainly boosted rate-distortion performance by refining two elements: the transform networks and the entropy model (Zafari et al. (2023); Mentzer et al. (2022); Ma et al. (2019); Lu et al. (2022); Liu et al. (2020); Gao et al. (2021); Fu et al. (2023); Bégaint et al. (2020); Li et al. (2025); Zhang et al. (2023)).
Early CNN variants introduced residual blocks (Cheng et al. (2020)) and octave residual modules (Chen et al. (2022)), while Xie et al. (2021) explored invertible architectures. Some works also employ deformable convolutions (Li et al. (2022a); Wang et al. (2023); Fu et al. (2024)) and GNNs (Tang et al. (2022); Yang et al. (2021); Spadaro et al. (2024) to achieve content-adaptive compression. Recently, Transformer-based architectures have surged in LIC research, including Swin-Transformer variants (Zou et al. (2022); Zhu et al. (2022)), hybrid CNN-Transformer schemes that fuse local detail with global context(Liu et al. (2023); Chen et al. (2025)), and frequency-aware window attention mechanisms (Li et al. (2024a)). Recently, researchers employ linear attention models, such as state space model Mamba (Gu and Dao (2023); Dao and Gu (2024)), for the non-linear transform networks to balance the RD performance and model complexity (Qin et al. (2024); Zeng et al. (2025)).
In terms of entropy models, advances include the autoregressive model by Minnen et al. (2018), checkerboard modeling (He et al. (2021)), channel-wise models (Minnen and Singh (2020)), and the space-channel context model of He et al. (2022). More works propose quadtree hierarchies (Li et al. (2022b, 2023)), Transformer-based entropy models (Qian et al. (2022)), and multi-reference schemes (Jiang et al. (2023); Jiang and Wang (2023)). T-CA further enhances channel-wise autoregression with transformers (Li et al. (2024a)). In this paper, we concentrate on refining the Mamba-based transform networks while leveraging the efficient SCCTX entropy model (He et al. (2022)).
2.2 State Space Models
State space models (SSMs) were first developed in control theory (Kalman (1960)). They are now popular again in deep learning because their sequential updates cost only linear time in the sequence length, which is ideal for modeling long-range relationships. S4 (Gu et al. (2021)) showed that deep SSMs can handle very long contexts, and S5 (Smith et al. (2022)) improved on this with a multi-input-multi-output SSM and a fast parallel-scan algorithm. Later, H3 (Fu et al. (2022)) reduced the performance gap between SSMs and Transformers on NLP tasks.
Mamba. The newest breakthrough is Mamba (Gu and Dao (2023); Dao and Gu (2024)). Mamba combines data-dependent gating with hardware-efficient kernels, reaching or exceeding Transformers while still scaling linearly. Researchers have already applied Mamba beyond language to vision tasks such as image classification (Zhu et al. ; Liu et al. (2024)), image segmentation (Yue et al. (2024)), and video understanding (Li et al. (2024b)). Some other works focus on low-level tasks: MambaIR (Guo et al. (2024b)) tackles local-pixel forgetting and channel redundancy of Mamba while applying it to single-image super-resolution; FreqMamba (Zhen et al. (2024)) operates in the Fourier domain for deraining; MambaLLIE (Weng et al. (2024)) modifies the state space equation for low-light enhancement; and further adaptations address dehazing, deblurring, rain removal, and fusion tasks (Zheng and Wu (2024); Gao et al. (2024); Wu et al. (2024); Xie et al. (2024); Qiao et al. (2024)). What’s more, SSMs are also adopted by learned image compression models such as Qin et al. (2024) and Zeng et al. (2025). Yet these works keep Mamba’s causal design, while 2-D image processing is naturally non-causal; this mismatch can shrink the effective receptive field and hurt RD performance.
3 Method
3.1 Preliminary
Fig. 2 (a) illustrates our learned image compression pipeline, which primarily comprises non-linear transform networks and an entropy model. The analysis transform $g_{a}(\,\cdot\,;\bm{\theta}_{a})$ encodes an input RGB image $\bm{x}$ into latent features $\bm{y}=g_{a}(\bm{x};\bm{\theta}_{a})$, where $\theta$ refers to the learned parameters. These features are discretized by uniform quantization, $\hat{\bm{y}}=Q(\bm{y}-\bm{\mu})+\bm{\mu}$, where the mean parameter $\bm{\mu}$ is provided by the entropy model. The synthesis transform $g_{s}(\,\cdot\,;\bm{\theta}_{s})$ then reconstructs the image $\hat{\bm{x}}=g_{s}(\hat{\bm{y}};\bm{\theta}_{s})$.
A Hyper-Encoder $h_{a}(\,\cdot\,;\bm{\phi}_{a})$ maps $\bm{y}$ to side-information $\bm{z}$, which is quantized to $\hat{\bm{z}}$. A Hyper-Decoder $h_{s}(\,\cdot\,;\bm{\phi}_{s})$ converts $\hat{\bm{z}}$ into Gaussian parameters $(\bm{\mu},\bm{\sigma})$. These parameters, together with spatial context $\bm{\varphi}$, define a Gaussian distribution $p_{\hat{\bm{y}}}\!\bigl{(}\hat{\bm{y}}\mid\bm{\mu},\bm{\sigma},\bm{\varphi}\bigr{)}$ used for entropy coding:
$\bm{z}=h_{a}(\bm{y};\bm{\phi}_{a}),\quad\hat{\bm{z}}=Q(\bm{z}),\quad\bm{\mu},\bm{\sigma}=h_{s}(\hat{\bm{z}};\bm{\phi}_{s}).$
$R=\mathbb{E}_{\bm{x}\sim p_{\bm{x}}}\!\bigl{[}-\!\log_{2}p_{\hat{\bm{y}}}(\hat{\bm{y}})\bigr{]}+\mathbb{E}_{\bm{x}\sim p_{\bm{x}}}\!\bigl{[}-\!\log_{2}p_{\hat{\bm{z}}}(\hat{\bm{z}})\bigr{]}.$
Figure 2: Overview of the Proposed Method. (a) The CMIC framework. Feature dimensions are set as {$C_{1},C_{2},C_{3},C_{4}$}, and the six non-linear transform stages have depths {$L_{1},L_{2},L_{3},L_{3},L_{2},L_{1}$}. Panel (b) shows the Content-Adaptive Mamba Block, detailing the Dynamic SSM architecture. Numbers 1-9 represent the indices of tokens, while letters a-d denote distinct cluster categories.
Figure 3: Our Entropy Model
Training objective. The network is optimized via the classic rate-distortion Lagrangian:
$\mathcal{L}_{\text{RD}}=R+\lambda\,\mathcal{D}(\bm{x},\hat{\bm{x}}),$
where $\mathcal{D}$ measures reconstruction error and the multiplier $\lambda$ tunes the bitrate-quality trade-off.
3.2 Overview of CMIC Model
First, we give a quick overview of how our Content-Adaptive Mamba-based LIC model (CMIC) works. As shown in Fig. 2 (a), CMIC follows a standard variational autoencoder layout. We split its non-linear transform into six stages based on feature resolution. Inspired by Liu et al. (2023); Li et al. (2024a), we utilize vanilla window-attention blocks to capture fine-grained local dependencies, while our proposed CAM blocks are introduced to enhance long-range modeling. A Conv-Feed Forward Network is added for efficient channel-wise information aggregation. Our entropy model is built upon the SCCTX entropy model (He et al. (2022)). As illustrated in Fig. 3, we incorporate depthwise convolution for efficient context modeling and employ a gated MLP for parameter aggregation. More details are provided in supplementary materials.
As shown in Fig. 2 (b), the core part of CAM block is a Dynamic SSM. It first clusters tokens into several categories, and reorders the flattened token sequence to bring similar tokens into proximity, enhancing local coherence in the token sequence. Meanwhile, instance-specific prompts are generated from the prompt dictionary. The reordered tokens, infused with these prompts, are scanned by a 1-D selective scan for efficient long-range modeling.
We detail two key components for Dynamic SSM: (i) Content-Adaptive Token Aggregation: A clustering mechanism groups and reorders tokens to encourage interaction between content-similar tokens. (ii) Global Priors Prompting: A learnable prompt dictionary generates instance-specific prompts with global priors to condition the selective scan.
3.3 Content-Adaptive Token Aggregation
Token Clustering.
Given the feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times d}$, the feature of each position is treated as a token $\mathbf{x}_{i}$, thus we have $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. We employed a cosine-based clustering to group tokens into $K$ clusters $\mathcal{G}_{1},\dots,\mathcal{G}_{K}$, each with normalized centers $\mathbf{c}_{1},\dots,\mathbf{c}_{K}$. To initialize the centers, the token sequence $\mathbf{X}$ is divided into $K$ consecutive segments, and each centroid $\mathbf{c}_{k}$ is computed as the average of the token embeddings within its corresponding segment. Each CAM block holds its own cluster centers, which are shared across all images and are only updated during training phase.
0: tokens $\mathbf{X}{=}\{\mathbf{x}_{i}\}_{1}^{N}$, centres $\mathbf{C}{=}\{\mathbf{c}_{j}\}_{1}^{K}$, iterations $T$, decay $\lambda$
1:for $t=1$ to $T$ do
2:  $\mathbf{D}_{i,j}\leftarrow\frac{\mathbf{x}_{i}^{\!\top}\mathbf{c}_{j}}{\lVert\mathbf{x}_{i}\rVert_{2}\,\lVert\mathbf{c}_{j}\rVert_{2}}$
3:  $g_{i}\leftarrow\arg\max_{j}\mathbf{D}_{i,j}$
4:  for $j=1$ to $K$ do
5:   if $\exists\,i:g_{i}{=}j$ then
6:    $\mathbf{c}^{\star}_{j}\leftarrow\frac{\sum_{i:g_{i}=j}\mathbf{x}_{i}}{\bigl{\lVert}\sum_{i:g_{i}=j}\mathbf{x}_{i}\bigr{\rVert}_{2}}$
7:   else
8:    $\mathbf{c}^{\star}_{j}\leftarrow\mathbf{c}_{j}$
9:   end if
10:  end for
11:end for
[…]
13:return  $\{g_{i}\}_{1}^{N},\mathbf{C}$
Algorithm 1 Training-phase Cosine K-Means
The K-Means algorithm (Algorithm 1) is employed in the training phase to update cluster centers for each training step. It begins by computing the cosine similarity between each token $x_{i}$ and every cluster center $c_{j}$, then assign each token to its nearest center, yielding assignments $\{g_{i}\}$. Next, for each cluster $j$, we compute the mean of all tokens assigned to it and  normalize it to obtain a tentative update $c_{j}^{*}$. If no tokens are assigned to cluster $j$, we keep its previous center unchanged to maintain stability. After $T$ iterations, we smooth the update via an exponential moving average (EMA) with decay $\lambda$. The algorithm outputs the final assignments $\{g_{i}\}$ (for sequence reordering) and the refined centers $\mathbf{C}$. During inference, we directly utilize the updated cluster centers to group the input tokens.
Sequence re-ordering.
Given the updated centroids, final assignments $g_{i}\in\{1,\dots,K\}$ is determined by the cosine similarity between tokens and cluster centers. We build a permutation $\pi:\{1,\dots,N\}\!\rightarrow\!\{1,\dots,N\}$ that first groups all tokens with $g_{i}=1$, then those with $g_{i}=2$, and so on. Applying $\pi$ yields a reordered sequence
$\tilde{\mathbf{X}}=\mathbf{X}_{\pi(\cdot)},$
so that tokens belonging to the same cluster become contiguous in 1-D sequence, enhancing local coherence of the sequence, and encouraging interactions between similar tokens. This permutation enables SSM to focus on proximity in feature space rather than in Euclidean space. The inverse permutation $\pi^{-1}$ is cached and applied to restore the original spatial layout.
3.4 Global Priors Prompting
Learnable Prompt Dictionary.
To enhance the model’s global awareness and mitigate the strict causality of Mamba, we introduce a learnable prompt dictionary $\mathbf{D}\in\mathbb{R}^{K\times d_{s}}$, where $K$ is the number of clusters and $d_{s}$ is the state space dimension of Mamba. Each prompt in this dictionary is a vector with size $d_{s}$ and corresponds to one of the clusters. As shown in Fig. 2 (b), the one-hot category assignment matrix $\bm{\Gamma}\in\{0,1\}^{N\times K}$ is formed by concatenating one-hot vectors, each representing the cluster membership of a token. The instance-specific prompts $\mathbf{P}\in\mathbb{R}^{N\times d_{s}}$ are generated via
$\mathbf{P}\;=\;\bm{\Gamma}\,\mathbf{D}.$
This dictionary learns typical prior knowledge from the training data, producing prompts that adapt to the input’s semantic layout. Combining global prototypes with instance-specific information, it enriches the model’s understanding of token relationships and global dependencies, effectively relaxing strict causality and improving long-range modeling.
Prompt Conditioning in Mamba.
The vanilla Mamba layer models token interactions with a discrete state-space system, expressed as:
$\bar{\mathbf{A}},\;\bar{\mathbf{B}}$ $\;=\;\exp(\Delta\mathbf{A}),\;\bigl{(}\Delta\mathbf{A}\bigr{)}^{-1}\!\!\bigl{(}\exp(\Delta\mathbf{A})-\mathbf{I}\bigr{)}\,\Delta\mathbf{B},$
$\mathbf{h}_{i}$ $\;=\;\bar{\mathbf{A}}\,\mathbf{h}_{i-1}+\bar{\mathbf{B}}\,\mathbf{x}_{i},$
$\mathbf{y}_{i}$ $\;=\;\mathbf{C}\,\mathbf{h}_{i}+\mathbf{D}\,\mathbf{x}_{i},$
where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is the $i$-th input token, $\mathbf{h}_{i}\in\mathbb{R}^{d_{s}}$ the hidden state, and $\mathbf{y}_{i}\in\mathbb{R}^{d}$ the output. Matrix $\bar{\mathbf{A}}$ acts as the state-transition or control matrix, $\bar{\mathbf{B}}$ as the input matrix, while $\mathbf{C}$ and $\mathbf{D}$ project the hidden state and the current input back to the output space. Following the Attentive State-Space Model of Guo et al. (2024a), we incorporate the global-prior prompt sequence $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d_{s}}$ by augmenting the output matrix:
$\mathbf{h}_{i}$ $=\bar{\mathbf{A}}\,\mathbf{h}_{i-1}+\bar{\mathbf{B}}\,\mathbf{x}_{i},$
$\mathbf{y}_{i}$ $=\bigl{(}\mathbf{C}+{\color[rgb]{0,0,1}\mathbf{P}}\bigr{)}\,\mathbf{h}_{i}+\mathbf{D}\,\mathbf{x}_{i},$
where $\mathbf{P}$ is the prompt vector encoding instance-specific global priors. The output kernel $\mathbf{C}$ functions analogously to the query in a self-attention layer Guo et al. (2024a): it determines which components of the hidden state are read out. By dynamically modulating the context matrix $\mathbf{C}$ with instance-specific prompts $\mathbf{P}$, the model is empowered to “query” information from unscanned tokens, effectively relaxing the strictly causal scan. Consequently, global priors that capture not only the instance-specific categorical information, but also tokens’ intricate categorical relationships are encoded in the dynamic SSM. This enhances SSM to reason about distant, unscanned tokens, enabling more content-adaptive long-range feature aggregation.
4 Experiments
4.1 Experimental Setup
[…]
4.2 Evaluation
We test our models on three image datasets: the Kodak dataset (Kodak (1993)) at 768×512768$\times$ 512768 × 512, the Tecnick test set (Asuni et al. (2014)) at $\times 1200$ 1200, and the CLIC professional validation set (CLIC (2021)) at 2K resolution. To assess rate-distortion performance, we report peak signal-to-noise ratio (PSNR), multiscale structural similarity (MS-SSIM), and bits per pixel (bpp). Average bitrate savings are quantified via BD-rate (Bjontegaard (2001)), and average PSNR improvement at matched bitrates via BD-PSNR. Model complexity is characterized by total parameter count, floating-point operations (FLOPs), peak GPU memory and decoding latency.
4.3 Rate-Distortion Performance
Our CMIC model is benchmarked against the conventional codec VTM-21.0 (Browne et al. ; Bross et al. (2021)), as well as several state-of-the-art (SOTA) learned methods including ELIC (He et al. (2022)), TCM (Liu et al. (2023)), MLIC++ (Jiang and Wang (2023)), FTIC (Li et al. (2024a)), CCA (Han et al. (2024)), and two Mamba-based LIC models, MambaVC (Qin et al. (2024)) and MambaIC (Zeng et al. (2025)). The resulting rate-distortion curves, shown in Fig. 56, demonstrate the performance trends across various bit rates. Comprehensive BD-rate comparisons, using VTM-21.0 as the reference, are detailed in Tab. 1. Our CMIC model achieves superior performance, reducing BD-rate by 15.91%, 21.34%, and 17.58% on the Kodak, Tecnick, and CLIC datasets, respectively.
Our experimental results demonstrate that the proposed CMIC model consistently outperforms competing methods across all datasets. Specifically, in terms of BD-PSNR improvements, CMIC achieves significant gains over the transformer-based LIC model FTIC, with improvements of 0.15 dB, 0.32 dB, and 0.36 dB on the Kodak, CLIC, and Tecnick datasets, respectively. Furthermore, the advantages become even more pronounced when compared to the CNN-transformer mixed model TCM-L, where CMIC achieves BD-PSNR increases of 0.28 dB, 0.40 dB, and 0.53 dB on the corresponding datasets. In terms of MS-SSIM, the CMIC model also achieves notable improvements, with gains of 0.52 dB and 0.30 dB on the CLIC dataset compared to TCM-L and FTIC, respectively, as shown in Fig. 6. These results not only highlight the robustness of our approach but also validate its superiority across a wide range of resolutions.
The CMIC model demonstrates a particularly strong performance on higher-resolution datasets, such as CLIC and Tecnick, underscoring the CAM’s effectiveness in global modeling. When compared to Mamba-based models, CMIC outperforms MambaVC by 7.51%, 6.80%, and 10.09% in BD-rate on the Kodak, CLIC, and Tecnick datasets, respectively. Similarly, CMIC surpasses MambaIC by 2.36%, 2.17%, and 6.48% in BD-rate on these datasets. As illustrated in Fig. 1 (c), CMIC consistently outperforms previous Mamba-based methods across the entire bitrate range.
Figure 4: RD Curves on the Kodak dataset.
Figure 5: RD Curves on the Tecnick dataset.
Figure 6: Performance evaluation on the CLIC Professional Validation Dataset.
  Method BD-rate (%) $\downarrow$ Complexity
Kodak Tecnick CLIC Params (M) FLOPs (T) Dec-Lat. (s) Peak-Mem (G)
VTM-21.0 0.00 0.00 0.00 - - - -
ELIC -3.10 -7.41 -0.84 33.29 1.74 0.335 1.50
MambaVC -8.10 -11.82 -10.94 47.88 2.10 0.425 14.73
CCA -11.99 -13.53 -11.40 64.89 3.28 0.385 5.04
FTIC -12.94 -13.89 -10.21 69.78 2.38 $>$10 4.90
MLIC++ -11.97 -15.13 -12.08 116.48 2.64 0.547 2.08
TCM -10.04 -10.42 -8.60 75.89 3.74 0.542 7.73
MambaIC -13.01 -15.27 -15.23 157.09 5.56 0.669 20.32
CMIC (Ours) -15.91 -21.34 -17.58 69.11 2.39 0.405 4.44
Table 1: Quantitative comparisons of SOTA LIC models. BD-rates are computed over VTM-21.0. FLOPs, peak memory and decoding latency are measured on 2K-resolution images. Since the official implementation of MambaIC is not released, we conducted a faithful reproduction based on the paper’s descriptions, achieving performance consistent with the original reported results.
4.4 Model Complexity
In addition to evaluating rate-distortion performance, we analyze model complexity by measuring parameter count, computational cost (FLOPs), decoding latency, and peak GPU memory consumption during inference on 2K resolution images (Tab. 1). Our proposed CMIC model achieves substantial BD-rate improvements on all three datasets, while maintaining moderate complexity levels (69.11M parameters, 2.39T FLOPs, 0.405s decoding latency, and 4.44GB peak memory). Compared to TCM, a previous SOTA LIC model with similar parameter scale (75.89M), CMIC exhibits significantly lower computational overhead in terms of FLOPs (36% lower), latency (25% faster), and GPU memory usage (43% less), despite delivering markedly superior compression performance. In comparison to the SOTA Mamba-based model, MambaIC, CMIC demonstrates clear advantages in parameter efficiency, FLOPs, and latency. Notably, it achieves a 57% reduction in GPU memory usage, a significant improvement attributed to its efficient single selective scan rather than the quadratic 2D scans This favorable complexity-performance balance is primarily attributed to our linear-complexity Mamba-based architecture, which efficiently captures global context without incurring excessive computational cost. These characteristics underscore CMIC’s practical advantages, making it particularly suitable for applications requiring high compression quality coupled with efficient processing.
4.5 Ablation Study and Analysis
  CTA GPP Kodak Tecnick CLIC
-13.26 -17.74 -14.87
-15.21 -20.17 -16.67
-14.27 -19.13 -15.34
-15.91 -21.34 -17.58
Table 2: Components ablation studies.
Content-Adaptive Token Aggregation. The concept of dynamic SSM is central to our approach. We first evaluate the contribution of Content-Adaptive Token Aggregation (CTA), implemented through K-means clustering and sequence re-ordering. As reported in Tab. 2, disabling both CTA and GPP means replacing CAM with vanilla Mamba. Enabling CTA yields BD-rate reductions of 1.7%-2.2% on Kodak, Tecnick, and CLIC, respectively. These gains confirm that scanning content-similar tokens consecutively—thus privileging proximity in feature space over Euclidean space—is crucial for the CAM block to capture long-range context.
Global Priors Prompting. To further adapt Mamba to image content, we inject instance-specific prompts, generated from a learnable dictionary, into the dynamic SSM. Tab. 2 shows that Global Priors Prompting (GPP) lowers BD-rate by 0.7%-1.2% on the three datasets. Together, those two techniques further boosts the BD-rate by -2.7%-3.6%, demonstrating that they are mutually supportive.
  BD-rate Params (M)
Conv Block -12.89 68.33
Swin Transformer -13.11 68.25
Window Attention -13.06 68.23
Standard Mamba -14.53 69.11
CAM (Ours) -15.91 69.11
Table 3: Ablations on network structures.
Comparison with Other Structures. To further validate the effectiveness of CAM blocks, we replace them with other basic blocks in stages 3-5 of the model. The results show that CAM blocks significantly outperform the alternatives while maintaining a comparable number of parameters. This highlights the efficiency and potential of CAM blocks for LIC models.
ERF Visualization. We visualize the Effective Receptive Field (ERF) (Luo et al. (2016)) of the analysis non-linear transforms in our CMIC model. We calculate the ERF as the magnitude of the gradient of the central latent pixel with respect to the input image, with values clipped to the range [0, 0.20]. The result is averaged over 24 images from the Kodak dataset. Compared with previous CNN-based, Transformer-based and Mamba-based LIC models, our CMIC model exhibits significantly larger receptive field across the whole image.
Figure 7: Effective Receptive Field (ERF) visualization for analysis networks in LIC models. Broader and darker green regions denote larger ERFs, which represents a more global spatial context.
Cluster Visualization. We further visualize the categorization results in Fig. 8 to verify the effectiveness of the clustering-based token aggregation mechanism. Each binary mask corresponds to a distinct cluster, and when compared with the original image, it becomes clear that tokens with similar visual or semantic attributes are grouped together (e.g., the red doors and windows in Kodim01, the clouds and sky in Kodim21, and the feathers in Kodim23). This demonstrates that content-similar tokens are reorganized into contiguous sequences, enabling the SSM to efficiently scan neighboring elements in the feature space rather than depending solely on Euclidean distance.
Figure 8: Visualization of clustering results. The examples presented are derived from Stage 2 of the analysis network. The clustering outcomes are visualized as binary (black-and-white) masks, where each panel represents a distinct cluster. Within a given panel, pixels rendered in white correspond to the associated cluster, while black pixels denote all other regions.
5 Limitation, Future Work and Broader Impact
Although our CMIC architecture achieves outstanding rate-distortion performance, it primarily focuses on aggregating content-similar tokens and fostering interactions among them. It does not enforce any specific scanning order across clusters or impose sorting mechanisms within individual clusters. In future work, we aim to introduce hierarchical reorganization and sorting strategies for both clusters and intra-cluster tokens, enhancing adaptability to diverse image content.
By exploring this dynamic SSM, our approach advances Mamba-based compression and fosters the evolution of adaptive LIC. It provides a foundation for refining Mamba models and developing adaptive strategies in the future. We hope this work inspires continued progress in these areas.
6 Conclusion
In this work, we introduced Content-Adaptive Mamba for Image Compression (CMIC), a novel approach tailored to align Mamba-style State Space models with the inherent two-dimensional structure of images. By employing clustering mechanism and a learnable prompt dictionary, we dynamically enhance the state-space model’s ability to capture long-range dependencies. Specifically, our content-adaptive scanning strategy sequentially processes tokens with similar content, emphasizing feature-space proximity rather than traditional Euclidean adjacency. Additionally, embedding global priors via the prompt dictionary relaxes the strict causality, significantly improving global context modeling. Consequently, CMIC efficiently aggregates global context while retaining linear computational complexity. Experimental results demonstrate that CMIC achieves SOTA RD performance, surpassing the traditional codec VTM-21.0 substantially across the Kodak, Tecnick, and CLIC datasets. These outcomes validate the effectiveness of our method in addressing the critical challenges of Mamba-based image compression.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly describe the claims and contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss about limitations in Sec. 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide all assumptions and correct proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all the information to reproduce our results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: We will release all the code and checkpoint on GitHub as soon as the paper is accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided all the details and would be happy to furnish any additional information that the reviewers, ACs, or PCs may request.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide the essential assumptions about the normal distribution. Error bars are not essential for our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide detailed information of computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We conform with all the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite and state all the correct essential sources.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.